- why does Ni + 4CO(g) = Ni(CO)4(g) reaction require T=350 K to occur and T=400 K for Ni(CO)4(g) to decompose?

- is this related to the splitting crystal field splitting ?

-what is the 18 rule of  Ni(CO)4(g) ?
Answer #1

The reaction between Ni and CO results in the production of Ni(CO)4. This is an inorganic complex involving a strong field ligand such as CO which is unsaturated thus allowing not only the ligand to σ−donate but also to accept electrons in its π* orbital from dπ metal orbitals and thereby making the CO ligand π−accepting. This makes the Nickel complex more stable and therefore requires more temperature to decompose.

The 18 rule is a rule in which if a molecule has an 18 electron count, it means that all of its orbitals are filled and it has achieved the noble gas type configuration.

In the complex, Ni(CO)4, Ni contributes 10 electrons (3d8 4s2) and ligand CO contributes 2 electrons. The total electrons become:-

Ni(CO)4 = 10 + 4(2) = 10 + 8 = 18 electrons

Don't forget to give a thumbs up if you think ur problem is resolved!!